OpenStudy (anonymous):

Tank A, measuring 91cm by 40cm by 41cm is half-filled with water. Tank B, measuring 40cm by 39cm by 88cm is empty. Without spilling from Tank A into Tank B till both water level were the same, what is the volume of water in Tank B now?

OpenStudy (anonymous):

I'm going to assume that the last number in each case is the height. The volume of water in each is going to be: \[91 \times 40 \times h \text{ for Tank A}\]\[40 \times 39 \times h \text{ for Tank B}\]

OpenStudy (luffingsails):

Hey BearBug, Sort of an involved problem. But I would approach the problem by writing down what I know. The important things to know is what is the volume of water that I have now and what are the equations for the water level of each tank. Water level of Tank A = (Water Volume of A) / (base of A x width of A) Water level of Tank B = (Water Volume of B) / (base of B x width of B) The total water volume is provided to us by your first statement: Water Volume = Water Volume of A + Water Volume of B Water Volume = (base of A) x (width of A) x (1/2)(height of A) Once you solve for the water volume you can then set the level equations equal and solve.

OpenStudy (luffingsails):

Sorry Dalv, didn't mean to cut you short. I was disconnected earlier.

OpenStudy (anonymous):

Yo, a good answer was provided in any case ;)